calculate the amount of electrical energy available from adry cell cooresponding to tje consumption of 6.55 g zinc metal(cell emf=1.5v, atm weight of zinc =65.5)


1
Expert's answer
2021-02-26T05:58:28-0500

n(Zn) = "\frac{6.55}{65.5}" = 0.1 (mol)

E = UQ = E(2nF) = 1.5(2"\times"0.1"\times"96485) = 28945.5 J.